OpenStudy (anonymous):

which equation represents the line passing through the points (4,-16) and is perpendicular to the line y=-2/3x+8? please help

OpenStudy (anonymous):

\[y=-\frac{ 2 }{ 3 }x+8\]

OpenStudy (mrdoe):

so to get the perpendicular line, you take the negative inverse of the slope in your first equation. so m=-2/3 ---> m=3/2 for the perpendicular line. so now just shift the line up or down until it passes through the point you want.

OpenStudy (zehanz):

The product of the slopes of perpendicular lines is -1. So if the first line has slope -2/3, the second has slope 3/2, because -2/3 * 3/2 = -1

OpenStudy (zehanz):

Once you have the slope, here 3/2, you fill in the coordinates of (4, -16) in the formula y= 3/2 *x + b. This will give you the number b.

OpenStudy (anonymous):

so the answer will be \[y=\frac{ 2 }{ 3 }x-22\]

OpenStudy (zehanz):

Right!